OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (lgbasallote):

\[\large \frac{x^2 + 2x +1}{x}\] first way to solve this: \[\large \frac{x^2}{x} + \frac{2x}{x} + \frac{1}{x}\] \[\large \implies x + 2 + \frac 1x\] second way to solve is by division x + 2 + 1/x ________________ x | x^2 + 2x + 1 - x^2 ------------- 2x - 2x --------- 1